OpenStudy (anonymous):

Solve for x: 7x + 15 = 38 twenty-three sevenths fifty-three sevenths 23 53

sammixboo (sammixboo):

First subtract 15 on both sides of the equation 7x + 15 - 15 = 38 - 15 7x = 38 - 15 38 - 15 is what?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

im sorry 15

sammixboo (sammixboo):

Not quite.

OpenStudy (onepieceftw):

23

sammixboo (sammixboo):

38 - 15 _____

sammixboo (sammixboo):

Please let the asker answer, thanks!

OpenStudy (anonymous):

sorry i looked at it wrong.

OpenStudy (anonymous):

and yes its 23..

sammixboo (sammixboo):

Right! So 7x = 23 Now you divide 7 on both sides of the equation 7x/7 = 23/7 x = 23/7 Can you simplify 23/7 for me since 7 doesn't go into 23 evenly

OpenStudy (anonymous):

i dont know how to simplify.

sammixboo (sammixboo):

Well how many times can 7 go into 23 evenly?

OpenStudy (anonymous):

none.

OpenStudy (anonymous):

it cant.

sammixboo (sammixboo):

We know that 7*3 = 21, but 7*4 = 28 So 7 can go into 23, 3 times but not perfectly

OpenStudy (anonymous):

yes

sammixboo (sammixboo):

Hold on wait

sammixboo (sammixboo):

We know that x = 23/7

sammixboo (sammixboo):

Look at the choices

OpenStudy (anonymous):

so the answer is 23?

OpenStudy (anonymous):

no no i mean 23/7

OpenStudy (anonymous):

no no i mean 23/7

sammixboo (sammixboo):

Yes

sammixboo (sammixboo):

:)